# On the Identifying Power of Generalized Monotonicity

### v3

@@ 3.2 Identifying Power of Generalized Monotonicity
Theorem 3.2 above establishes that, for possibly multi-valued $Y$, $D$, and $Z$, and any model $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ such that $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$, if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, then $\Theta_{0}(P,\mathbf{Q})$ equals (5). We now show that $\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ as long as $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$, so that the identified set for $\theta(Q)$ assuming instrument exogeneity and generalized monotonicity coincides with the identified set assuming instrument exogeneity alone, as long as both assumptions are consistent with the distribution of the observed data. Our result therefore generalizes Balke and Pearl (1993, 1997), which study the case of binary $Y$, $Z$, and $D$.
In order to do so, we consider a mean independence assumption even weaker than Assumption 2.1, and show the identified set under this even weaker assumption is also (5). A sandwich argument will then lead to our desired result. In particular, we first establish that the identified set for $\theta(Q)$ in (5) coincides with the identified set under the weaker mean independence assumption considered by Robins (1989) and Manski (1990):
@@
[…]
@@
[…]
@@
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$. Then, the sets in (5) and (8) coincide.
Using Lemma 3.2, we are able to establish our desired result, which asserts that, maintaining Assumption 2.1, additionally imposing Assumption 2.2 either causes the identified set for $\theta(Q)$ to become empty (if those assumptions are not consistent with $P$) or leaves the identified set for $\theta(Q)$ unchanged (if those assumptions are consistent with $P$).
@@
An implication of Theorem 3.3 is that the identified set on $\theta(Q)$ under Assumption 2.1 alone will be (5) whether or not one imposes Assumption 2.2 as long as the distribution of the data is consistent with Assumptions 2.1 and 2.2.
@@
We now consider some restrictions on potential treatments that have been considered previously in the literature. In each case, we show $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$; in particular, these restrictions satisfy generalized monotonicity. We emphasize that frequently $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$. In what follows, it is implicitly understood that Assumption 3.1 is satisfied, so that the model imposes no restriction on potential outcomes. Thus, in each example Theorem 3.1 applies and such restrictions do not provide any identifying power for average potential outcomes or ATEs. In Appendix C, we consider three additional examples: an RCT with a “close substitute” as considered in Kline and Walters (2016); the monotonicity and “irrelevance” assumptions considered in Kirkeboen et al. (2016); and an additive random utility model for a binary treatment.
@@
Consider a multi-arm randomized controlled trial (RCT) with noncompliance, where $Z=d$ denotes random assignment to treatment $d$, $D_{d}=d$ denotes that the subject would comply with assignment if assigned to treatment $d$, and $|\mathcal{D}|=|\mathcal{Z}|$. More generally, not necessarily in the context of an RCT, one can interpret $Z=d$ as encouragement to treatment $d$ and interpret $D_{d}=d$ as the subject would take treatment $d$ if encouraged to do so. In this example, $Q$ satisfies Assumption 2.1 because $Z$ is randomly assigned. We may generalize the “no-defier” restriction of Angrist et al. (1996) as: for each $d\in\mathcal{D}$,
@@
Cheng and Small (2006) consider an RCT with noncompliance where $\mathcal{D}=\mathcal{Z}=\{0,1,2\}$. Assumption 2.1 continues to hold because $Z$ is randomly assigned. In such a setting, they develop bounds on average effects within subgroups defined by potential treatments which, following the terminology of Frangakis and Rubin (2002), they call “principal strata.” While Bai et al. (2024) derives the identified sets for those parameters given their assumptions, we now use our analysis to consider instead identification of average potential outcomes and ATEs given their assumptions. Their “Monotonicity I” assumption is equivalent to one-sided noncompliance in the preceding example. Their “Monotonicity II” assumption states that subjects who would comply with assignment to treatment $2$ would also comply with assignment to treatment $1$, so that
@@
Suppose $Q$ satisfies Assumption 2.1. Heckman and Pinto (2018) define “unordered monotonicity” as the assumption that, for any $d\in\mathcal{D}$, and any $z,z^{\prime}\in\mathcal{Z}$,
$Q\{\mathbbm{1}\{D_{z}=d\}\geq\mathbbm{1}\{D_{z^{\prime}}=d\}\}=1\text{ or }Q\{\mathbbm{1}\{D_{z}=d\}\leq\mathbbm{1}\{D_{z^{\prime}}=d\}\}=1~{}.$ (9)
@@
The ARUM for a binary treatment is equivalent to the Heckman-Vytlacil nonparametric selection model for a binary treatment considered, e.g., in Heckman and Vytlacil (1999, 2005), and shown by Vytlacil (2002) to be equivalent to the monotonicity and exogeneity assumptions of Imbens and Angrist (1994). Heckman and Vytlacil (2001) show that the Heckman-Vytlacil nonparametric selection model for a binary treatment results in an identified set for the ATE of the form in (6) and that the model has no identifying power beyond instrument exogeneity for the ATE. Example C.1 in Appendix C shows that an ARUM for a binary treatment will satisfy Assumption 2.2 and therefore the results of this paper nest the results of Heckman and Vytlacil (2001). Example 4.4, on the other hand, extends their results to nonparametric selection models for a multi-valued treatment. For a partial identification analysis of a class of parameters that includes the ATE under a nonparametric selection model for a binary treatment and sometimes imposing additional restrictions, see, e.g., Mogstad et al. (2018), Han and Yang (2024), and Marx (2024).
@@
Lee and Salanié (2023) also consider the ARUM defined by (10) without imposing the uniform targeting of (11) for each treatment. Instead, they impose an assumption that they refer to as “strict one-to-one targeting”, in which the set of treatments can be partitioned into a set of treatments $\mathcal{D}^{\dagger}$ that are “targeted” and a set of treatments $\mathcal{D}\setminus\mathcal{D}^{\dagger}$ that are “not targeted” such that
@@
    For $d\in\mathcal{D}^{\dagger}$, there exists $z^{\dagger}(d)$ such that $g(z^{\dagger}(d),d)>g(z^{\prime},d)$ for all $z^{\prime}\neq z^{\dagger}(d)$ and such that $g(z^{\prime},d)$ takes the same value for all $z^{\prime}\neq z^{\dagger}(d)$; additionally, $z^{\dagger}(d)\neq z^{\dagger}(d^{\prime})$ for $d,d^{\prime}\in\mathcal{D}^{\dagger}$, $d\neq d^{\prime}$.
The terminology “one-to-one” stems from the second requirement above. They further impose that there exists a treatment that is known to be non-targeted. This class of ARUMs is equivalent to imposing (12) for targeted treatments, imposing $g(z,d)=\alpha_{d}$ for non-targeted treatments, and imposing that there is at least one non-targeted treatment. In such a setting, Lee and Salanié (2023) analyze the identification of a particular class of average effects within subgroups defined by potential treatments. We now use our analysis to consider instead the identification of average potential outcomes and ATEs given their assumptions. Suppose strict one-to-one targeting holds, and additionally suppose that there is at least one targeted treatment. In Appendix B.1, we argue that Assumption 2.2 holds when $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$, so that there are more values of the instrument than there are targeted treatments. We further argue that (11) does not hold for some treatments unless $|\mathcal{D}|=|\mathcal{Z}|=2$. Such models therefore provide another class of ARUMs, distinct from the one with uniform targeting described in Example 4.4, for which Assumption 2.2 holds. In Example 5.3 below, we show, however, that Assumption 2.2 does not hold when $|\mathcal{D}|\geq 3$, $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|$, and the support of $(U_{d}:d\in\mathcal{D})$ is $\Re^{|\mathcal{D}|}$.
@@
which, in the language of Angrist et al. (1996), is imposing that all individuals are either compliers or defiers. For any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, $\theta(Q)$ is identified relative to $\mathbf{Q}$, i.e., $\Theta_{0}(P,\mathbf{Q})$ is a singleton. To see this, note that for any $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$ and $y\in\mathcal{Y}$,
@@
For example, $D$ might represent quantity of some treatment, and $Z$ might represent levels of subsidy for the treatment. The restriction in (13) is equivalent to the monotonicity assumption considered in Angrist and Imbens (1995). See Vytlacil (2006) for the connection between this restriction and ordered discrete-choice selection models. Without loss of generality, let $\mathcal{D}=\{0,\ldots,\bar{D}\}$. In this case, (3) is satisfied for $d\in\{0,\bar{D}\}$ for all $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$; to see this, take $z^{\ast}(0)=\min\{\mathcal{Z}\}$ and $z^{\ast}(\bar{D})=\max\{\mathcal{Z}\}$. By a straightforward modification of the arguments underlying Theorem 3.2, one can show that the identified sets for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$ are given by (5) for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. The ordered monotonicity assumption in (13) therefore has no identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$. In contrast, (3) need not hold for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$ and $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. In Appendix B.3, we show there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We further show $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ for the same $P$. Thus, the ordered monotonicity assumption in (13) does have identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{d}]$ for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$.
@@
In Example 4.5, we considered ARUMs satisfying the strict one-to-one targeting assumption of Lee and Salanié (2023). As discussed there, if $|\mathcal{D}|=2$ or if $|\mathcal{D}|\geq 3$ and $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$, then Assumption 2.2 holds and the results in Section 3 are applicable. Now consider the case in which $|\mathcal{D}|\geq 3$ and $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|$. Denote by $\mathbf{Q}$ the ARUM model defined by (10) under the additional assumption that the support of $(U_{d}:d\in\mathcal{D})$ is $\Re^{|\mathcal{D}|}$. Then, as we show in Appendix B.4, while (3) will hold for targeted treatments, (3) cannot hold for any non-targeted treatment, and thus Assumption 2.2 is violated. We show that, while the identified set for $\mathbb{E}_{Q}[Y_{d}]$ is given by (5) for targeted treatments when $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and the identified set for $\mathbb{E}_{Q}[Y_{d}]$ is not given by (5) for non-targeted treatments.

### v4

@@ 3.2 Identifying Power of Generalized Monotonicity
Theorem 3.2 above establishes that, for possibly multi-valued $Y$, $D$, and $Z$, and any model $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ such that $\mathbf{Q}$ does not restrict the potential outcomes, if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, then $\Theta_{0}(P,\mathbf{Q})$ equals (6). We now show that $\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ as long as $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$, so that the identified set for $\theta(Q)$ assuming instrument exogeneity and generalized monotonicity coincides with the identified set assuming instrument exogeneity alone, as long as both assumptions are consistent with the distribution of the observed data. Our result therefore generalizes balke1993nonparametric; balke1997bounds, which study the case of binary $Y$, $Z$, and $D$.
In order to do so, we consider a mean independence assumption even weaker than Assumption 2.1, and show the identified set under this even weaker assumption is also (6). A sandwich argument will then lead to our desired result. In particular, we first establish that the identified set for $\theta(Q)$ in (6) coincides with the identified set under the weaker mean independence assumption considered by robins1989analysis and manski1990nonparametric:
@@
Note Assumption 3.2 is weaker than instrument exogeneity in Assumption 2.1, and does not imply (2). Let $\mathbf{Q}_{\it MI}^{\ast}$ denote the set of all $Q$ that satisfies Assumption 3.2 (where $MI$ stands for mean independence). Following robins1989analysis and manski1990nonparametric, the following lemma derives the identified set for $\theta(Q)$ under mean independence:
@@
Suppose $\mathbf{Q}_{0}(P,\mathbf{Q}_{\it MI}^{\ast})\neq\emptyset$. Then,
$\Theta_{0}(P,\mathbf{Q}_{\it MI}^{\ast})=\prod_{d\in\mathcal{D}}\left[\max_{z\in\mathcal{Z}}\{\beta_{d|z}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd\mid z})\},~\min_{z\in\mathcal{Z}}\{\beta_{d|z}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd\mid z})\}\right].$ (9)
@@
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$, $\mathbf{Q}$ satisfies Assumption 3.1, and $P$ is such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Then, the sets in (6) and (9) coincide.
Using Lemma 3.2, we are able to establish our desired result, which asserts that, maintaining Assumption 2.1, additionally imposing Assumption 2.2 either causes the identified set for $\theta(Q)$ to become empty (if those assumptions are not consistent with $P$) or leaves the identified set for $\theta(Q)$ unchanged (if those assumptions are consistent with $P$). In fact, we will establish a stronger result, that if a submodel of instrument exogeneity and generalized monotonicity is consistent with $P$, then any model sandwiched between this submodel and the model that only assumes mean independence leads to the same identified set for $\theta(Q)$.
@@
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}$ satisfies Assumption 3.1. Further suppose $\mathbf{Q}^{\prime}$ satisfies
$\mathbf{Q}\subseteq\mathbf{Q}^{\prime}\subseteq\mathbf{Q}_{\it MI}^{\ast}~.$
Then, for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$,
$\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}^{\prime})=\Theta_{0}(P,\mathbf{Q}_{\it MI}^{\ast})~.$
Remark 3.3.
Theorem 3.3 implies that in order for a model to have identifying power for average potential outcomes, it has to be the case that the model does not contain a submodel of instrument exogeneity and generalized monotonicity that is consistent with $P$. In other words, the model has to contradict Assumption 2.1 or 2.2. We illustrate this observation in Example 5.1 below.
Corollary 3.2.
@@
An implication of Corollary 3.2 is that the identified set for $\theta(Q)$ under Assumption 2.1 alone will be (6), regardless of whether Assumption 2.2 is imposed, as long as the distribution of the data is consistent with Assumptions 2.1 and 2.2.
Theorem 3.3 further implies that under the following weaker exogeneity assumption, generalized monotonicity also has no identifying power for average potential outcomes and ATEs:
Assumption 3.3 (Weak Instrument Exogeneity).
See richardson2013single for an analysis of alternative exogeneity restrictions, and in particular how Assumption 3.3 avoids the “cross-world” restrictions of the stronger joint independence in Assumption 2.1. Denote by $\mathbf{Q}_{\it WE}^{\ast}$ the set of all distributions that satisfy Assumption 3.3 and $\mathbf{Q}_{\it WE,M}^{\ast}$ the set of all distributions that satisfy Assumptions 3.3 and 2.2.
Corollary 3.3.
For any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q}_{\it WE,M}^{\ast})\neq\emptyset$, $\Theta_{0}(P,\mathbf{Q}_{\it WE,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{\it WE}^{\ast})$.
@@
We now consider some restrictions on potential treatments that have been considered previously in the literature. In each case, we show $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$; in particular, these restrictions satisfy generalized monotonicity. We emphasize that frequently $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$. In what follows, it is implicitly understood that Assumption 3.1 is satisfied, so that the model imposes no restriction on potential outcomes. Thus, in each example Theorem 3.1 applies and such restrictions do not provide any identifying power for average potential outcomes or ATEs. In Appendix C, we consider three additional examples: an RCT with a “close substitute” as considered in kline2016evaluating; the monotonicity and “irrelevance” assumptions considered in kirkeboen2016field; and an additive random utility model for a binary treatment.
@@
Consider a multi-arm randomized controlled trial (RCT) with noncompliance, where $Z=d$ denotes random assignment to treatment $d$, $D_{d}=d$ denotes that the subject would comply with assignment if assigned to treatment $d$, and $|\mathcal{D}|=|\mathcal{Z}|$. More generally, not necessarily in the context of an RCT, one can interpret $Z=d$ as encouragement to treatment $d$ and interpret $D_{d}=d$ as the subject would take treatment $d$ if encouraged to do so. In this example, $Q$ satisfies Assumption 2.1 because $Z$ is randomly assigned. We may generalize the “no-defier” restriction of angrist1996identification as: for each $d\in\mathcal{D}$,
@@
cheng2006bounds consider an RCT with noncompliance where $\mathcal{D}=\mathcal{Z}=\{0,1,2\}$. Assumption 2.1 continues to hold because $Z$ is randomly assigned. In such a setting, they develop bounds on average effects within subgroups defined by potential treatments which, following the terminology of frangakis2002principal, they call “principal strata.” While bai2025inference derives the identified sets for those parameters given their assumptions, we now use our analysis to consider instead identification of average potential outcomes and ATEs given their assumptions. Their “Monotonicity I” assumption is equivalent to one-sided noncompliance in the preceding example. Their “Monotonicity II” assumption states that subjects who would comply with assignment to treatment $2$ would also comply with assignment to treatment $1$, so that
@@
Suppose $Q$ satisfies Assumption 2.1. heckman2018unordered define “unordered monotonicity” as the assumption that, for any $d\in\mathcal{D}$, and any $z,z^{\prime}\in\mathcal{Z}$,
[…]
@@
The ARUM for a binary treatment is equivalent to the Heckman-Vytlacil nonparametric selection model for a binary treatment considered, e.g., in heckman1999local; heckman2005structural, and shown by vytlacil2002independence to be equivalent to the monotonicity and exogeneity assumptions of imbens1994identification. heckman2001instrumental show that the Heckman-Vytlacil nonparametric selection model for a binary treatment results in an identified set for the ATE of the form in (7) and that the model has no identifying power beyond instrument exogeneity for the ATE. Example C.1 in Appendix C shows that an ARUM for a binary treatment will satisfy Assumption 2.2 and therefore the results of this paper nest the results of heckman2001instrumental. Example 4.4, on the other hand, extends their results to nonparametric selection models for a multi-valued treatment. For a partial identification analysis of a class of parameters that includes the ATE under a nonparametric selection model for a binary treatment and sometimes imposing additional restrictions, see, e.g., mogstad2018using, han2024computational, and marx2024sharp.
@@
lee2023treatment also consider the ARUM defined by (11) without imposing the uniform targeting of (12) for each treatment. Instead, they impose an assumption that they refer to as “strict one-to-one targeting”, in which the set of treatments can be partitioned into a set of treatments $\mathcal{D}^{\dagger}$ that are “targeted” and a set of treatments $\mathcal{D}\setminus\mathcal{D}^{\dagger}$ that are “not targeted” such that
@@
    For $d\in\mathcal{D}^{\dagger}$, there exists $z^{\dagger}(d)$ such that $g(z^{\dagger}(d),d)>g(z^{\prime},d)$ for all $z^{\prime}\neq z^{\dagger}(d)$ and such that $g(z^{\prime},d)$ takes the same value for all $z^{\prime}\neq z^{\dagger}(d)$; additionally, $z^{\dagger}(d)\neq z^{\dagger}(d^{\prime})$ for $^{\prime}\in\mathcal{D}^{\dagger}, ddd\neq d^{\prime}$.
The terminology “one-to-one” stems from the second requirement above. They further impose that there exists a treatment that is known to be non-targeted. This class of ARUMs is equivalent to imposing (13) for targeted treatments, imposing $g(z,d)=\alpha_{d}$ for non-targeted treatments, and imposing that there is at least one non-targeted treatment. In such a setting, lee2023treatment analyze the identification of a particular class of average effects within subgroups defined by potential treatments. We now use our analysis to consider instead the identification of average potential outcomes and ATEs given their assumptions. Suppose strict one-to-one targeting holds, and additionally suppose that there is at least one targeted treatment. In Appendix B.1, we argue that Assumption 2.2 holds when $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$, so that there are more values of the instrument than there are targeted treatments. We further argue that (12) does not hold for some treatments unless $|\mathcal{D}|=|\mathcal{Z}|=2$. Such models therefore provide another class of ARUMs, distinct from the one with uniform targeting described in Example 4.4, for which Assumption 2.2 holds. In Example 5.3 below, we show, however, that Assumption 2.2 does not hold when $|\mathcal{D}|\geq 3$, $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|$, and the support of $(U_{d}:d\in\mathcal{D})$ is $\Re^{|\mathcal{D}|}$.
@@
which, in the language of angrist1996identification, is imposing that all individuals are either compliers or defiers. For any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, $\theta(Q)$ is identified relative to $\mathbf{Q}$, i.e., $\Theta_{0}(P,\mathbf{Q})$ is a singleton. To see this, note that for any $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$ and $y\in\mathcal{Y}$,
@@
Furthermore, recall as discussed in Remark 3.3 that if $\mathbf{Q}$ has identifying power for $\theta(Q)$, then it cannot contain a submodel satisfying instrument exogeneity and generalized monotonicity that is consistent with $P$. In this example, the model $\mathbf{Q}$ contains such a submodel if and only if $P$ satisfies either (i) $P\{D=0\mid Z=0\}=P\{D=1\mid Z=1\}=1$ (in which case all individuals are compliers) or (ii) $P\{D=1\mid Z=0\}=P\{D=0\mid Z=1\}=1$ (in which case all individuals are defiers). In order for $\mathbf{Q}$ to have identifying power for $\theta(Q)$, it therefore must be the case that $0<P\{D=0\mid Z=0\},P\{D=0\mid Z=1\}<1$, which is indeed satisfied by the counterexample in Appendix B.2. Further note that (6) is a singleton in this example if and only if $P$ satisfies either (i) or (ii).
@@
For example, $D$ might represent quantity of some treatment, and $Z$ might represent levels of subsidy for the treatment. The restriction in (14) is equivalent to the monotonicity assumption considered in angrist1995two. See vytlacil2006ordered for the connection between this restriction and ordered discrete-choice selection models. Without loss of generality, let $\mathcal{D}=\{0,\ldots,\bar{D}\}$. In this case, (4) is satisfied for $d\in\{0,\bar{D}\}$ for all $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$; to see this, take $z^{\ast}(0)=\min\{\mathcal{Z}\}$ and $z^{\ast}(\bar{D})=\max\{\mathcal{Z}\}$. By a straightforward modification of the arguments underlying Theorem 3.2, one can show that the identified sets for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$ are given by (6) for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. The ordered monotonicity assumption in (14) therefore has no identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$. In contrast, (4) need not hold for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$ and $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. In Appendix B.3, we show there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and $\Theta_{0}(P,\mathbf{Q})$ is not given by (6). We further show $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ for the same $P$. Thus, the ordered monotonicity assumption in (14) does have identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{d}]$ for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$.
@@
In Example 4.5, we considered ARUMs satisfying the strict one-to-one targeting assumption of lee2023treatment. As discussed there, if $|\mathcal{D}|=2$ or if $|\mathcal{D}|\geq 3$ and $|\mathcal{Z}|>|\mathcal{D}^{\dagger}|$, then Assumption 2.2 holds and the results in Section 3 are applicable. Now consider the case in which $|\mathcal{D}|\geq 3$ and $|\mathcal{Z}|=|\mathcal{D}^{\dagger}|$. Denote by $\mathbf{Q}$ the ARUM model defined by (11) under the additional assumption that the support of $(U_{d}:d\in\mathcal{D})$ is $\Re^{|\mathcal{D}|}$. Then, as we show in Appendix B.4, while (4) will hold for targeted treatments, (4) cannot hold for any non-targeted treatment, and thus Assumption 2.2 is violated. We show that, while the identified set for $\mathbb{E}_{Q}[Y_{d}]$ is given by (6) for targeted treatments when $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and the identified set for $\mathbb{E}_{Q}[Y_{d}]$ is not given by (6) for non-targeted treatments.
Disclosure Statement
The authors report there are no competing interests to declare.